What could NOT be the sides of a right triangle?
1. (3,4,5)
2. (13,15,17)
3. (9,12,15)
4. (8,15,17)

Answer by jim_thompson5910(35256)

The answer is choice 2 because


13^2 + 15^2 = 169 + 225 = 394 which is NOT equal to 17^2 (17^2 = 289)